# Nonequilibrium physics of brain dynamics

### v1

@@ 3.1.4 Neural field theory
[…]
@@ 3.3.2 Machine learning the arrow of time
Another framework for measuring the level of nonequilibrium in time-series data is to use a machine learning approach [230]. To do this, the problem is recast as a guessing game. If a machine learning algorithm can correctly classify forward trajectories of a system, $\mathbf{x}_{f}(t)$, and backward (time-reversed) trajectories, $\mathbf{x}_{b}(t)$, then it has learnt the arrow of time. More specifically, we consider a thermodynamic system defined by a Hamiltonian where $W$ is the work performed on the system during a trajectory and $F=U-TS$ is the Helmholtz free energy, where $U$ is the internal energy, $T$ is the temperature and $S$ is the entropy [231]. The change in free energy between the initial and final states is $\Delta F=F_{B}-F_{A}$. The probability that the trajectory was generated by the forward process is,
@@ 3.3.3 Time-series symbolisation
A central challenge for calculating information-theoretic quantities in neuroimaging, is that MVTS evolve in a continuous state-space with infinite possible values (up to the resolution of the measurement). One classic approach to solving this problem is to transform a continuous MVTS, into a univariate series of discrete symbols [233]. A very common symbolisation method was proposed by Bandt and Pompe for the calculation of permutation entropy [234], which can be applied to univariate trajectories. Given a series $x_{t}\in\mathbb{R}$ for $t=1,...,T$, one considers a time-delay embedding into a higher dimension $D$ [235],
@@
Each time-point in the embedded series $\bm{y}_{t}$ is assigned a ordinal pattern which is a permutation of $(1,...,D)$ representing the order of the values in the embedded state-vector at that time-point. As a result, each time-point is now represented by a pattern $\bm{z}_{t}=\{p_{i,t},i\in[1,...,D]\}$, where the pattern is one of $D!$ finite possible patterns or symbols. Panels A & B of Fig. 7 illustrates ordinal pattern analysis applied to a stochastic trajectory and its time-reversal. Such a symbolisation can be used to measure the time-irreversibility of the observed trajectory [49]. Given a series $x_{t}$ and its time reversal $\hat{x}_{t}$, one calculates the series of ordinal patterns for each i.e. $\bm{z}_{t}$ and $\hat{\bm{z}}_{t}$ respectively. Next, one calculates the discrete probability distributions, $P$ and $\hat{P}$, which measure the probability of observing each ordinal pattern in both the forward and backward trajectory. The KL divergence between $P$ and $\hat{P}$ becomes a natural measure of the irreversibility of the process [49].
@@ 3.3.4 Visibility graphs
Visibility graphs (VG) encode temporal correlations in their degree distribution [237]. Given a time-series $\{x_{t}\}_{t=1,...,T}$, the VG has one node for each $t=1,...,T$. Nodes $i$ and $j$ are connected by an edge if the corresponding points $(i,x_{i})$ and $(j,x_{j})$ are ‘mutually visible’ which means that they satisfy that, for any intermediate data-point $(k,x_{k})$, with $i<k<j$,
@@ 3.4.1 Irreversibility and higher-order interactions
Figure 8: Higher-order interactions from irreversible neural dynamics. A. Higher-order interactions represent a generalisation of networks to relationships between groups of more than two variables. Higher-order interactions can either be represented by hypergraphs or simplicial complexes, where simplicial complexes necessarily require that all lower-order interactions are included to defined a higher-order interaction e.g. the existence of a triangle requires the existence of all its edges. Black lines represent pairwise edges, whilst pink triangles represent 3-order hyperedges. B. A schematic representation of 5 brain regions in the auditory system, where one aims to infer which higher-order interactions are particularly irreversible. Processing regions are denoted in red with sensory regions in blue. The regions are auditory cortices on the left and right (ACL, ACR), median cingulate gyrus (MC), ventro-medial prefrontal cortex (VMPFC) and hippocampal regions on the left and right (HITL, HITR). Adapted from Ref. [52] C. The irreversibility of pairwise and triplet interactions shows which pairs/triplet have a particularly irreversible interaction as measured by the DiMViGI framework. The number of (*$\dagger$) represent the number of standard deviations the irreversibility is away from the mean across all pairs/triplets. The icon along the $x$-axis denotes each pairwise and triplet interaction in the system with black circles denoting included regions, with reference to the schematic in Panel B. Adapted from Ref. [52].
@@ 3.4.2 Information decomposition of irreversibility
This decomposition, known as ‘$\Phi$ID’, captures the way in which information is transformed over time within a pairwise interaction. Upon time-reversal, the total TDMI remains constant between $t$ and $t^{\prime}$, yet the information dynamics are reversed. As a result, $\Phi$ID provides a family of metrics to quantify irreversibilities arising from three different kinds of information dynamics [249]. Although the applications of such a decomposition have yet to be explored in empirical recordings of neural activity, it is known that cross-correlation measures are insufficient to capture some components of temporal irreversibility. Furthermore, such measures have been used to analyse simulated mean-field dynamics of neural populations revealing synergistic contributions to the overall level of irreversibility [249]. This approach represents a promising avenue for gaining further insight into the information dynamics underlying nonequilibrium brain dynamics.
@@ 3.4.3 Decomposition of linear dynamics into oscillatory modes
where $\chi_{k}=\lambda_{k}/2\pi\mathbf{i}$ and $\lambda_{k}$ is the $k-$th eigenvalue of $-\bm{B}$, the friction matrix, appearing in Eq. (23). Here we have defined the following,
@@ 4 Analysis of neural spike-trains
In order to relate spiking dynamics to a discrete-state model, one starts with a raster plot, as shown in Panel A of Fig. 10, which records the exact spike times of each neuron in the recording. By considering a sliding window of a given length, one converts a raster plot into a discrete-time process where a neuron is active, in a given window, if it has at least one spike, and inactive otherwise. This coding yields a binary spike matrix, as in Panel B of Fig. 10, which can be considered to be an observation of a stochastic process in discrete-time with states in the space $\{0,1\}^{N}$, where $N$ is the number of neurons [^6] Or {±$\{\pm 1\}^{N}$ in the case of the Ising model..
@@ 4.1 The asymmetric kinetic Ising model
Figure 10: Neural spike trains and the inverse Ising problem. A. An example of a neural spike-train raster plot. For each of three neurons, the exact times of spikes are recorded. By considering a sliding window of a certain length, this can be converted into a binary spike matrix. B. A binary spike matrix represents units that are firing in a given window with a ’1’ and units that are not firing with a ’0’. C. A spike matrix for a population of 67 neurons. This output is comparable to the dynamics of an Ising model. D. Using gradient descent or mean-field methods, we can infer an asymmetric network of interactions $bm{J}$ where $J_{ij}$ represents the coupling strength from neuron $j$ to $i$. E. In addition, we infer the external input $H_{i}$ to each neuron $i$.
@@ 4.2 Decomposing irreversibility in spike-trains
where $P(s_{i},s_{i}^{\prime};s_{j})$ represents the joint transition probability of variable $i$ given that variable $j$ remains fixed at $s_{j}$. By choosing the marginal distribution, $P_{i}(s_{i},s_{i}^{\prime};\bm{s}_{-i})$, that minimises (140) and satisfies (144), one calculates the minimum EPR that is consistent with observed pairwise dynamics, which is denoted $\Phi^{(2)}$. In general, one can then define $\Phi^{(k)}$, for $k=1,...,N$ where $\Phi^{(1)}=\Phi^{\text{ind}}$ and $\Phi^{(N)}=\Phi$ [53]. Finally, the unique component of the EPR that emerges from interactions at order $k$ can be defined recursively to be,
@@ 4.3 Self-organised criticality in neural activity
One of the most popular avenues for applying statistical physics techniques to neural systems is the study of criticality. In statistical physics, criticality describes the end-point of a phase transition, which is typically a discontinuous change in a medium in response to external conditions [278]. One example is the equilibrium Ising model on a 2D lattice with nearest-neighbour coupling, in the absense of external fields, where the system is operating at temperature $T$. For $T$ less than the critical value $T_{c}$, the system is magnetised, meaning the vast majority of the spins have the same state, either ±$\pm 1$, also known as ferromagnetic long-range order. For temperatures higher than $T_{c}$, the system is globally disordered, meaning there are no clusters of magnetised spins. At the critical point $T=T_{c}$, the correlation length, which measures the spatial distance over which two spins are correlated, diverges to infinity [278]. The correlation length is the length scale associated with the exponential decay of the correlation function:
@@
[…]
@@
In the context of neural dynamics, much of the focus has been on the emergence of avalanches [286, 287]. In order to assess if the brain is operating at, or near, the critical point, the cascade of neural activity that follows a neuron firing can be described as an avalanche whose size [^7] This includes the duration and the number of neurons involved can be estimated from data. If the size of avalanches in the brain follows a power-law distribution, it can be seen as evidence of SOC [286], as avalanches in critical sand-pile models follow a similar distribution [281]. The critical brain hypothesis argues that self-organising to criticality can optimise the speed of information processing in the brain due to high responsiveness to external stimulus without the risk of signals failing to propagate, nor triggering runaway excitation [286]. Moreover, some systems admit a Griffith’s phase, which is an entire region of critical-like dynamics [288], as opposed to a single point, which have since been shown to emerge in neuronal dynamics [289].
@@ 6 Discussion and outlook
Approaches integrating dynamical systems, statistical physics, and complex systems have generated many new innovative methods for analysing neural system recordings [305, 306, 3, 307]. Although a relatively recent perspective in neuroscience, the study of nonequilibrium dynamics in biological systems has been essential for understanding their mechanisms and energetics [11, 12], particularly through its connections to stochastic thermodynamics [308]. However, interpretation of nonequilibrium dynamics from the perspective of thermodynamics is less obvious for the brain since the relationship between meso/macroscopic recorded signals and microscopic energetic processes becomes blurred across scales. Nevertheless, the formalism of stochastic nonequilibrium dynamics has repeatedly proved fruitful in the analysis of neural recordings. For instance, accumulating evidence increasingly suggests that neural dynamics can be well approximated by a stochastic process in a nonequilibrium steady state [25, 30, 54] and that the entropy production rate is a robust neural correlate of consciousness [28, 42, 40]. Unlike traditional neural correlates of consciousness [309] which originate from neuroscience, nonequilibrium measures come equipped with the formal mathematical apparatus developed for stochastic thermodynamics, which allows for different, but complementary, formulations of the same phenomena. An example of this is the nonequilibrium theories of discrete Markov chains and continuous Langevin processes, presented in Sec. 2, both of which have been applied to the study of nonequilibrium brain dynamics from two inherently different perspectives, yet they yield consistent results.
@@
Nonequilibrium brain dynamics is the latest attempt to leverage tools from the quiver of dynamical systems and statistical physics for the analysis of neural dynamics across time- and length-scales. It has lead to the development of important correlates of cognitive exertion and consciousness that offer a bridge between dynamics and neuro-psychology. Beyond merely quantifying the global irreversibility of neural signals, this approach has driven the development of innovative modelling and analysis techniques that emphasize asymmetry and hierarchy in brain organization. While still in its early stages, this line of research holds significant promise for advancing our understanding of neural dynamics through the lens of nonequilibrium physics, opening new avenues for both theory and application in neuroscience.

### v2

@@ 3.1.4 Neural field theory
$\frac{1}{\alpha}\partial_{t}u(x,t)$ $=-u(x,t)+\int_{-\infty}^{\infty}dy\;w(y)f\circ u(x-y,t),$ (89)
@@ 3.2.2 The Gaussian hidden Markov model
Table 1 summarises the model-based approaches to nonequilibrium neural dynamics.
Model Ref. Dynamics Topology Data Type
Linear Langevin Sec. 3.1.1 Continuous Network Continuous
Kuramoto Sec. 3.1.2 Continuous Network Phase time-series
Hopf Sec. 3.1.3 Continuous Network Continuous/Oscillatory
Neural field theory Sec. 3.1.4 Continuous Geometric mesh Continuous
General Langevin Sec. 3.1.5 Continuous Network Continuous
Clustered Markov chain Sec. 3.2.1 Discrete - Continuous
Hidden Markov model Sec. 3.2.2 Discrete - Continuous
Ising Sec. 4.1 Discrete Network Spike-trains
Table 1: Model-based approaches to nonequilibrium neural dynamics. This summarises the model-based approaches described in this paper and which mathematical framework they fall under.
@@ 3.3.2 Machine learning the arrow of time
Another framework for measuring the level of nonequilibrium in time-series data is to use a machine learning approach [243]. To do this, the problem is recast as a guessing game. If a machine learning algorithm can correctly classify forward trajectories of a system, $\mathbf{x}_{f}(t)$, and backward (time-reversed) trajectories, $\mathbf{x}_{b}(t)$, then it has learnt the arrow of time. More specifically, we consider a thermodynamic system defined by a Hamiltonian where $W$ is the work performed on the system during a trajectory and F=UTSF=U-TS is the Helmholtz free energy, where $U$ is the internal energy, $T$ is the temperature and $S$ is the entropy [244]. The change in free energy between the initial and final states is $\Delta F=F_{B}-F_{A}$. The probability that the trajectory was generated by the forward process is,
@@ 3.3.3 Time-series symbolisation
A central challenge for calculating information-theoretic quantities in neuroimaging, is that MVTS evolve in a continuous state-space with infinite possible values (up to the resolution of the measurement). One classic approach to solving this problem is to transform a continuous MVTS, into a univariate series of discrete symbols [246]. A very common symbolisation method was proposed by Bandt and Pompe for the calculation of permutation entropy [247], which can be applied to univariate trajectories. Given a series $x_{t}\in\mathbb{R}$ for t=1,,Tt=1,...,T, one considers a time-delay embedding into a higher dimension $D$ [248],
@@
Each time-point in the embedded series $\bm{y}_{t}$ is assigned a ordinal pattern which is a permutation of (1,,D)(1,...,D) representing the order of the values in the embedded state-vector at that time-point. As a result, each time-point is now represented by a pattern $\bm{z}_{t}=\{p_{i,t},i\in[1,...,D]\}$, where the pattern is one of $D!$ finite possible patterns or symbols. Panels A & B of Fig. 8 illustrates ordinal pattern analysis applied to a stochastic trajectory and its time-reversal. Such a symbolisation can be used to measure the time-irreversibility of the observed trajectory [53]. Given a series $x_{t}$ and its time reversal $\hat{x}_{t}$, one calculates the series of ordinal patterns for each i.e. $\bm{z}_{t}$ and $\hat{\bm{z}}_{t}$ respectively. Next, one calculates the discrete probability distributions, $P$ and $\hat{P}$, which measure the probability of observing each ordinal pattern in both the forward and backward trajectory. The KL divergence between $P$ and $\hat{P}$ becomes a natural measure of the irreversibility of the process [53].
@@ 3.3.4 Visibility graphs
Visibility graphs (VG) encode temporal correlations in their degree distribution [250]. Given a time-series $\{x_{t}\}_{t=1,...,T}$, the VG has one node for each t=1,,Tt=1,...,T. Nodes $i$ and $j$ are connected by an edge if the corresponding points $(i,x_{i})$ and $(j,x_{j})$ are ‘mutually visible’ which means that they satisfy that, for any intermediate data-point $(k,x_{k})$, with $i<k<j$,
@@ 3.4.1 Irreversibility and higher-order interactions
Figure 9: Higher-order interactions from irreversible neural dynamics. A. Higher-order interactions represent a generalisation of networks to relationships between groups of more than two variables. Higher-order interactions can either be represented by hypergraphs or simplicial complexes, where simplicial complexes necessarily require that all lower-order interactions are included to defined a higher-order interaction e.g. the existence of a triangle requires the existence of all its edges. Black lines represent pairwise edges, whilst pink triangles represent 3-order hyperedges. B. A schematic representation of 5 brain regions in the auditory system, where one aims to infer which higher-order interactions are particularly irreversible. Processing regions are denoted in red with sensory regions in blue. The regions are auditory cortices on the left and right (ACL, ACR), median cingulate gyrus (MC), ventro-medial prefrontal cortex (VMPFC) and hippocampal regions on the left and right (HITL, HITR). Adapted from Ref. [56] C. The irreversibility of pairwise and triplet interactions shows which pairs/triplet have a particularly irreversible interaction as measured by the DiMViGI framework. The number of ($\dagger$) represent the number of standard deviations the irreversibility is away from the mean across all pairs/triplets. The icon along the $x$-axis denotes each pairwise and triplet interaction in the system with black circles denoting included regions, with reference to the schematic in Panel B. Adapted from Ref. [56].
@@ 3.4.2 Information decomposition of irreversibility
[…]
@@ 3.4.3 Decomposition of linear dynamics into oscillatory modes
where $\chi_{k}=\lambda_{k}/2\pi\mathbf{i}$ and $\lambda_{k}$ is the $k$-th eigenvalue of $\bm{B}$, the friction matrix, appearing in Eq. (23). Here we have defined the following,
@@ 3.5.3 Testing for stationarity in neural data
3.5.4 Preprocessing neural data
Neural data from all modalities are contaminated by noise and artifacts. When analysing neural data, a consistent and principled preprocessing pipeline is crucial to obtain valid results. Best practices are continually evolving within the distinct communities of fMRI, MEG, EEG and electrophysiology. Here, we will briefly comment on preprocessing approaches and their possible modulation of irreversibility metrics. However, there is no correct answer on how to best preprocess data, and selecting between preprocessing options is more an art form than a hard science and left up to the neuroscientists’ best judgement.
Preprocessing stages can alter both the spatial and temporal information in a signal. For the purposes of measuring irreversibility, it is the preservation of temporal structure that is most important. For example, one may opt to use a causal filter, which performs filtering using only previous time-steps rather than future information, implicitly respecting the arrow of time [278]. Beyond opting for causal methods when possible, we recommend following widely tested and accepted approaches such as correcting for spatial and gradient distortions, head motion, intensity normalization, and bias field removal (in fMRI) - for example through the HCP minimal preprocessing pipeline [279, 280]. Moreover, structural artifacts can be removed using independent component analysis approaches [281]. A number of open-source toolboxes exist for performing such preprocessing in a reproducible manner such as FMRIB Software Library (FSL), FreeSurfer, and FieldTrip [282].
When analysing spike train data, such as in Sec. 4, it is typical to have a prohibitively fine temporal granularity, thus data must be temporally coarse-grained in order to apply computational methods in a realistic time-frame. It is known that temporal coarse-graining masks irreversibility, leading to underestimates in the EPR [37]. Accounting for this coarse-graining can be performed via finite data correction as described in Sec. 3.5.2. When combining a reproducible, principled and minimalistic preprocessing pipeline with rigorous surrogate data testing, one can be more assured that results pertaining to the irreversibility of neural signals stem from genuine nonequilibrium dynamics, rather than artifacts and noise, but the possible effect of preprocessing should never be ignored.
@@ 4 Analysis of neural spike-trains
In order to relate spiking dynamics to a discrete-state model, one starts with a raster plot, as shown in Panel A of Fig. 11, which records the exact spike times of each neuron in the recording. By considering a sliding window of a given length, one converts a raster plot into a discrete-time process where a neuron is active, in a given window, if it has at least one spike, and inactive otherwise. This coding yields a binary spike matrix, as in Panel B of Fig. 11, which can be considered to be an observation of a stochastic process in discrete-time with states in the space $\{0,1\}^{N}$, where $N$ is the number of neurons [^7] Or $\{\pm 1\}^{N}$ in the case of the Ising model..
@@ 4.1 The asymmetric kinetic Ising model
Figure 11: Neural spike trains and the inverse Ising problem. A. An example of a neural spike-train raster plot. For each of three neurons, the exact times of spikes are recorded. By considering a sliding window of a certain length, this can be converted into a binary spike matrix. B. A binary spike matrix represents units that are firing in a given window with a ’1’ and units that are not firing with a ’0’. C. A spike matrix for a population of 67 neurons. This output is comparable to the dynamics of an Ising model. D. Using gradient descent or mean-field methods, we can infer an asymmetric network of interactions bmJbm{J} where $J_{ij}$ represents the coupling strength from neuron $j$ to ii. E. In addition, we infer the external input $H_{i}$ to each neuron ii.
@@ 4.2 Decomposing irreversibility in spike-trains
where $P(s_{i},s_{i}^{\prime};s_{j})$ represents the joint transition probability of variable $i$ given that variable $j$ remains fixed at $s_{j}$. By choosing the marginal distribution, $P_{i}(s_{i},s_{i}^{\prime};\bm{s}_{-i})$, that minimises (140) and satisfies (144), one calculates the minimum EPR that is consistent with observed pairwise dynamics, which is denoted $\Phi^{(2)}$. In general, one can then define $\Phi^{(k)}$, for k=1,,Nk=1,...,N where $\Phi^{(1)}=\Phi^{\text{ind}}$ and $\Phi^{(N)}=\Phi$ [57]. Finally, the unique component of the EPR that emerges from interactions at order $k$ can be defined recursively to be,
@@ 4.3 Self-organised criticality in neural activity
One of the most popular avenues for applying statistical physics techniques to neural systems is the study of criticality. In statistical physics, criticality describes the end-point of a phase transition, which is typically a discontinuous change in a medium in response to external conditions [296]. One example is the equilibrium Ising model on a 2D lattice with nearest-neighbour coupling, in the absense of external fields, where the system is operating at temperature $T$. For $T$ less than the critical value $T_{c}$, the system is magnetised, meaning the vast majority of the spins have the same state, either $\pm 1$, also known as ferromagnetic long-range order. For temperatures higher than $T_{c}$, the system is globally disordered, meaning there are no clusters of magnetised spins. At the critical point $T=T_{c}$, the correlation length, which measures the spatial distance over which two spins are correlated, diverges to infinity [296]. The correlation length is the length scale associated with the exponential decay of the correlation function:
@@
where $r=d(i,j)$ is the distance between spins $i$ and $j$ and $\langle\cdot\rangle$ is the average. Above the critical temperature, this function decays exponentially,
@@
In the context of neural dynamics, much of the focus has been on the emergence of avalanches [304, 305]. In order to assess if the brain is operating at, or near, the critical point, the cascade of neural activity that follows a neuron firing can be described as an avalanche whose size [^8] This includes the duration and the number of neurons involved can be estimated from data. If the size of avalanches in the brain follows a power-law distribution, it can be seen as evidence of SOC [304], as avalanches in critical sand-pile models follow a similar distribution [299]. The critical brain hypothesis argues that self-organising to criticality can optimise the speed of information processing in the brain due to high responsiveness to external stimulus without the risk of signals failing to propagate, nor triggering runaway excitation [304]. Moreover, some systems admit a Griffith’s phase, which is an entire region of critical-like dynamics [306], as opposed to a single point. Griffith’s phases have since been shown to emerge in neuronal dynamics [307].
@@ 6 Discussion and outlook
Approaches integrating nonlinear dynamics, statistical physics, and complex systems have generated many new innovative methods for analysing neural system recordings [323, 324, 3, 325]. Although a relatively recent perspective in neuroscience, the study of nonequilibrium dynamics in biological systems has been essential for understanding their mechanisms and energetics [11, 12], particularly through its connections to stochastic thermodynamics [326]. However, interpretation of nonequilibrium dynamics from the perspective of thermodynamics is less obvious for the brain since the relationship between meso/macroscopic recorded signals and microscopic energetic processes becomes blurred across scales. Nevertheless, the formalism of stochastic nonequilibrium dynamics has repeatedly proved fruitful in the analysis of neural recordings. For instance, accumulating evidence increasingly suggests that neural dynamics can be well approximated by a stochastic process in a nonequilibrium steady state [25, 30, 58] and that the entropy production rate is a robust neural correlate of consciousness [28, 46, 44]. Unlike traditional neural correlates of consciousness [60] which originate from neuroscience, nonequilibrium measures come equipped with the formal mathematical apparatus developed for stochastic thermodynamics, which allows for different, but complementary, formulations of the same phenomena. An example of this is the nonequilibrium theories of discrete Markov chains and continuous Langevin processes, presented in Sec. 2, both of which have been applied to the study of nonequilibrium brain dynamics from two inherently different perspectives, yet they yield consistent results.
@@
A central question in neural dynamics, that has not yet been discussed here, is how they perform distributed computation. Notably, it is believed that computation occurs at the neuronal level, thus it is more opaque to methods focusing on large-scale brain dynamics. However, as nonequilibrium dynamics emerge across scales, their role in computation may be significant. For example, as mentioned previously, Hopfield’s model of associative memory relies on symmetric coupling and reversible dynamics to encode memories in a potential landscape [77, 333]. Nonequilibrium currents, enforced via asymmetric coupling [79], could be used to design models of associative memory with improved capacity [334] or to induce cycling between a sequence of memories. Whilst this was explored in the wake of Hopfield’s original papers [335, 336, 337], the focus was on random asymmetric couplings, leading to limited biological relevance and interest. Modern research on nonequilibrium neural systems could rejuvenate this dormant area, leading to nonequilibrium models of associative memory.
More modern research on computation via neural dynamics focuses on the emergence of neural manifolds which are low-dimensional manifold structures present in high-dimensional neural recordings [338]. Analysing dynamics over the manifold [339], as well as symmetry breaking [340], from the perspective of nonequilibrium dynamics could bring new insight into how neural systems compute through active dynamics. The relationship between neural systems and artificial recurrent neural networks also provides an avenue to study machine learning systems with nonequilibrium dynamics. Somewhat aside from neuroscience, applications of stochastic thermodynamics have emerged in deep learning such as through the study of stochastic optimisation algorithms [341] or through new architectures like diffusion models [342].
Nonequilibrium brain dynamics is the latest attempt to leverage tools from the quiver of dynamical systems and statistical physics for the analysis of neural dynamics across time- and length-scales. It has lead to the development of important correlates of cognitive exertion and consciousness that offer a bridge between dynamics and neuro-psychology. Beyond merely quantifying the global irreversibility of neural signals, this approach has driven the development of innovative modelling and analysis techniques that emphasize asymmetry and hierarchy in brain organization.
Whilst theoretical, mathematical frameworks have lead to significant developments in physics, they have so far failed to answer questions of equal fundamentality in neuroscience and biology. By introducing the techniques of nonequilibrium statistical physics, we make progress towards a more complete theoretical understanding of the brain within a paradigm that holds the potential to draw links between the brain as a dynamical system, biological organ and information processing machine. Nevertheless, this field is in its infancy, and more work is needed to develop this area into a unified framework that is theoretically consistent and empirically useful.